OpenStudy (anonymous):

m<1 and m<2 are a linear pair. m<1= x- 35, and m<2= x+ 83. Find the measure of each angle

OpenStudy (anonymous):

Linear pair => Sum of both angles is 180 degrees. So, x - 35 + x + 83 = 180 => 2x = 132 => x = 66 Angle 1 = 66 - 35 = 31 Angle 2 = 66 + 83 = 149